OpenStudy (jjuden):

divide assume the denominator does not equal zero (18x^5y^3 + 6x^2y^4 + 24x^6y) ÷ 3xy

OpenStudy (campbell_st):

my advise it to split it into smaller problems \[\frac{18x^5y^3}{3xy} + \frac{6x^2y^4}{3xy} + \frac{24x^6y}{3xy}\] then approach it term by term the 1st fraction can be simplified to \[6x^4y^2 + \frac{6x^2y^4}{3xy} + \frac{24x^6y}{3xy}\] you'll find 3xy divides evenly into each